OpenStudy (anonymous):

@Mathmatics Ashley spent one and one-half hours on math homework, one-fourth hour on music homework, 25 minutes on language arts homework, and one-half of an hour on her science fair project. How much total time did she spend on her homework? 2 hours 2 hours and 30 minutes 2 hours and 40 minutes 2 hours and 45 minutes

OpenStudy (anonymous):

im stuck!!!!!!!!

OpenStudy (vishal_kothari):

2 hours and 40 minutes....

OpenStudy (anonymous):

how

OpenStudy (vishal_kothari):

1.5+0.25+0.416+0.5

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

90min+15min+25min+30min=160mins 160-60-60=40 answer = 2 hours 40 minutes

OpenStudy (vishal_kothari):

just add it , you got the answer....

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{1}{1} + \frac{1}{2} + \frac{1}{4} + \frac{{25}}{{60}} + \frac{1}{2}\\ = \frac{{60}}{{60}} + \frac{{30}}{{60}} + \frac{{15}}{{60}} + \frac{{25}}{{60}} + \frac{{30}}{{60}}\\ = \frac{{60 + 30 + 15 + 25 + 30}}{{60}}\\ = \frac{{160}}{{60}}\\ = \frac{8}{3}\\ = 2 + \frac{2}{3} \end{array}\]

OpenStudy (anonymous):

2 hours and 40. minutes i think

OpenStudy (anonymous):

thanks everybody wow!

OpenStudy (anonymous):

2 hours + (2/3)*60 = 2 hours + 40 minutes.

OpenStudy (moongazer):

2 hours and 40 minutes